When f(x) = 3x² + 2x +ă find the rate of change by finding hange by mindig f(b)-f(a) b-a f(6) - f(a) b - a Preview
When f(x) = 3x² + 6x + 2, find the rate of change by finding f(0) - f(a) f(6) - f(a) b - a P review
h(b) – h(a) When h(t) = 2t + 3t - 6, find the rate of change by finding – 2 When b=9 and a=3 h(9) – h(3) 9-3
When P(t) = 1t2 - 1t - 3, find the rate of change by finding P(6) - P(a) w When b=7 and a=4 b-a P(7) - P(4) 7- 4
Answer #1

Q-1 f(x) = 382 + 2x + 7 f(b) - fia) b-a - (362 +26+7) - (392+29+7) 16-9) = 3662-93) +26b-a) to (b-a) = 3(b-q) (bta) + 2(6-9)Q.3. (t) = 2+2 +37-6 a bcb) -hca) - (262+36-6) -(2q2 +39-6) (6-9) (b-a) - 266²-02) + 3(b-a) to (b-a0 - zob-a) (bta) +3(6-9) (Q.4. Plt) = 1.t? - Lt - 3 - PCD) - P(a) =(B2-6-3) - (QP-2-3) (b-a) (b-a) 0 + (079-)+(26-29) = (b-a) = (b-a)(b+a) + 60cb-a) C

Thank You !☺️